PLEASE HELP!

​​​​​​​

  1. Lane works in Quebec and is paid on a semi-monthly basis. He has the following types of earnings & benefits. Determine which are pensionable and insurable and then calculate the CPP and EI deductions.
  1. Statutory Holiday Pay $120
  2. Regular $1200
  3. Employer paid Parking $30
Answer #1

Pensionable earnings composed of persons income from an office which are salary, wages or any other remuneration including tips and gratuities received by that person in a year. So regular pay, Saturday holiday pay and employer paid parking forms pensionable earnings.

Insurable earning does not include non cash benefits. Since employer paid parking is non cash benefit it is not included in Insurable earnings.

Calculation of CPP

Total pensionable income = $1200 + $120 + $ 30 = $ 1350. (Step 1)

Basic pay period exemption for 2020 (Semi- mostly basis) =   $145.83 (Step 2)

Step 1 - Step 2 = $1350 - $ 145.83 = $ 1204.17 (Step 3)   

Step 3 * CPP contribution rate 2020. = $1204.17 * 5.25 %. =. $ 63.2189

So CPP for the period = $ 63.2189

Calculation of EL

Insurable Amount = $ 1200 + $ 120 = $ 1320 (Step 1)

Employee EL premium rate for 2020 = 1.58 % (Step 2)

EL for the period = 1320 * 1.58% = $ 20.856